Find a polynomial f(x) of degree 3 with real coefficients and the following zeros.
-1,-1+i

Answer by stanbon(75887)

Find a polynomial f(x) of degree 3 with real coefficients and the following zeros.
-1,-1+i
---
If the coefficients are Real, -1-i must also be a zero.
----
f(x) = (x+1)(x-(-1+i))(x-(-1-i))
---
f(x) = (x+1)((x+1)-i)((x+1)+i))
---
f(x) = (x+1)((x+1)^2 + 1)
-----
f(x) = (x+1)(x^2+2x+2)
---
f(x) = x^3 +2x^2+x^2 + 2x+2x + 2
---
f(x) = x^3+4x^2+4x+2
----------------
Cheers,
Stan H.
---------------